write the equation of a circle with center (2,4) and containing the point (-2,1)
Answer by lwsshak3(11628)

write the equation of a circle with center (2,4) and containing the point (-2,1)
**
Standard form of equation of a circle: (x-h)^2+(y-k)^2=r^2, (h,k)=(x,y) coordinates of center, r=radius
For given circle:
(x-2)^2+(y-4)^2=r^2
plug in (x,y) coordinates of given point (-2,1) on the circle to find r^2
(-2-2)^2+(1-4)^2=r^2
(-4)^2+3^2=16+9=25=r^2
Equation of given circle:
(x-2)^2+(y-4)^2=25